OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. cosine of pi divided by five times cosine of pi divided by seven plus sine of pi divided by five times sine of pi divided by seven.

OpenStudy (arthur_ser):

\[\frac{\frac{\frac{\frac{ \cos \pi }{ 5\cos \pi} }{ 7+\sin \pi }}{ 5\sin \pi} }{ 7}\] like that?

OpenStudy (anonymous):

\[\cos \pi/5 \times \cos \pi/7 + \sin \pi/5 \times \sin \pi/7\]

OpenStudy (anonymous):

@arthur_ser

OpenStudy (arthur_ser):

\[=\cos \left( \frac{ \pi }{ 5 }-\frac{ \pi }{ 7 } \right)=\cos \left( \frac{ 7\pi }{ 35 }-\frac{ 5\pi }{ 35 } \right)=\cos \left( \frac{ 2\pi }{ 35} \right)\] I think..